# Strength Lies in Differences! Towards Effective Non-collaborative Dialogues via Tailored Strategy Planning

### v2

@@ Abstract
Strength Lies in Differences! Towards Effective Non-collaborative Dialogues via Tailored Strategy Planning
@@ 1 Introduction
Non-collaborative dialogues, such as negotiation He et al. (2018) and persuasion Wang et al. (2019), occur when the agent and user hold conflicting interests Deng et al. (2023b, a). Typically, both parties need to employ various strategies to achieve an agreement favorable to themselves Keizer et al. (2017); Zhan et al. (2024). As user resistance varies depending on the agent’s strategies Shi et al. (2019); Dutt et al. (2021), it is imperative for the agent to perform strategic planning tailored to diverse users. Relying on a one-size-fits-all strategy can leave the agent vulnerable to others taking advantage due to its lack of adaptability and flexibility Yang et al. (2021); Xu et al. (2023).
Recent efforts have resorted to large language models (LLMs) as dialogue agents to perform non-collaborative tasks Deng et al. (2023b); Fu et al. (2023); Zhang et al. (2023a). They aim to guide the response of LLMs through mixed-initiative prompts Chen et al. (2023); Deng et al. (2023b); Zhang et al. (2023a) or incorporating an external strategy planner Yu et al. (2023); Deng et al. (2023c). However, these initiatives has been criticized regarding its performance in real-world scenarios Deng et al. (2023c); Kwon et al. (2024), where users have various non-collaborative strategies. We attribute this outcome to the neglect of two crucial aspects: 1) Existing methods fail to incorporate explicit user-specific characteristics into their strategic planning, instead relying solely on the conversational history. Importantly, by creating informative representations of individual users, agents can adapt their behaviors and devise tailored strategies Jang et al. (2020); Yang et al. (2021). 2) Their training paradigm fails to generate strategic planners that generalize well to diverse users. Their paradigms are oversimplified, relying on a single user simulator for interactive training. This simulator is restricted in generating varied non-collaborative behaviors, often exhibiting a focus on prioritizing user contentment Zhang et al. (2023c); Durmus et al. (2023); Bianchi et al. (2024). Essentially, agents trained in this manner are accustomed to engage with a single user exclusively, leading to rigidity and obstinacy when encountering new users with different interaction behaviors Wang et al. (2023); Safdari et al. (2023).
@@
To tackle these challenges, we design a simple yet effective method, called Trip, to improve LLMs’ capability in Tailored stRategIc Planning. Trip includes a user-aware strategic planning module and a population-based training paradigm. Specifically, the strategic planning module incorporates user-specific characteristics into strategic planning using the Theory-of-Mind (ToM) Premack and Woodruff (1978); Wimmer and Perner (1983). This involves analyzing users’ mental states and future possible actions during interactions to understand their interests Yang et al. (2021); Chawla et al. (2023a). Moreover, instead of relying on a solitary user simulator, our population-based training paradigm promotes the adaptation of the strategic planning module to various users, achieved by training it with more diverse user simulators. Each simulator is equipped with extensive sets of non-collaborative strategies and role-playing personas. As such, Trip essentially manipulates the experience of the dialogue agent, enabling it to recognize the importance of tailoring strategies for individual users. Our key contributions are concluded below:
@@ 2 Related Work
Strategic planning for non-collaborative dialogues. Recent researches have introduced various methods based on LLMs to enhance their effectiveness in strategic planning. These methods can be categorized into two types: 1) Developing stimulus prompts to unleash the potential of LLMs. Chen et al. (2023) validate the effectiveness of using mixed-initiative prompts to tackle proactive dialogue challenges. Deng et al. (2023b) and Zhang et al. (2023a) encourage LLMs to engage in self-reflection to plan their next actions. Fu et al. (2023) employ self-play simulations to iteratively refine strategic planning by soliciting feedback from other LLMs. Nonetheless, as highlighted by Deng et al. (2023c), the effectiveness of these approaches is impeded by non-trainable parameters. 2) Equipping LLMs with an external strategy planner. The planner is capable of generating prompts at each turn, providing nuanced, instance-specific guidance and control over LLMs. This could be integrated using methods like Monte Carlo Tree Search Yu et al. (2023) or a plug-in model Deng et al. (2023c), which can be fine-tuned for improving the strategic planning capability without affecting the functionalities of LLM-powered dialogue agents. However, these methods still struggle to achieve promising results due to their inability to integrate user-specific characteristics into their strategic planning. Complementary to Deng et al. (2023c), our work investigates the importance of tailored strategic planning by modeling user-related characteristics explicitly.
Training paradigms for non-collaborative dialogues. Current training paradigms involve the dialogue agent interacting with a single user simulator to enhance its strategic planning capabilities. In specific, Chawla et al. (2023b) build a user simulator that mimics human-human dialogue data in a supervised manner, while Yu et al. (2023); Deng et al. (2023c) resort to a role-playing LLM-based user simulator. However, a single user simulator can only represent the behaviors of one or a type of users, potentially leading to the under-representation of other users’ behaviors, as evidenced by Liu et al. (2023); Shi et al. (2019). Therefore, existing training paradigms fail to produce strategic planners that cater to diverse users with varying behaviors. In this paper, our work investigates the importance of tailored strategic planning by diversifying the user’s behaviors using population-based training.
@@ 3.1 Evaluation Setup
Evaluation Overview. The environment encompasses various synthetic user simulators showcasing diverse non-collaborative behaviors. In the evaluation process, each dialogue agent must interact with these simulators Deng et al. (2023c). During their interactions, the dialogue agent and user simulator alternate in employing strategies in their responses with the ultimate aim of maximizing their own self-interest. The interactions continues until the conversational goal is achieved or the maximum number of turns is reached. We gather these interactions and assess the agents performances.
@@
Baselines. We consider two representative baselines: Standard agent (i.e., vanilla LLM without any modification) and PPDPP agent Deng et al. (2023c), which is current SOTA agent with a trainable external strategy planner [^1] Notably, we also consider other existing dialogue agents in our main experiments..
@@
Diverse User Simulators. Our simulators are synthesized with non-collaborative behaviors, guided by their task-relevant personas. As evidenced by previous study Deng et al. (2023a); Bianchi et al. (2024), LLMs are limited to demonstrate non-collaborative behaviors. To this end, we prompt non-collaborative behaviors explicitly into LLMs using the resisting strategies that are designed to foil persuasion attempts Fransen et al. (2015); Tian et al. (2020); Dutt et al. (2021). Initially, we equip LLMs with different personas Jiang et al. (2023); Zhou et al. (2023b); Zhang et al. (2023b), which are used to select non-collaborative behaviors from the set of resisting strategies. Following Wang et al. (2019); Jiang et al. (2024), we consider two types of personas, including Big-Five Personality [^2] Openness, Conscientiousness, Extraversion, Agreeableness, and Neuroticism Goldberg (1992) and Decision-Making Styles [^3] Directive, Conceptual, Analytical, and Behavioral Scott and Bruce (1995), together with LLM-generated cohesive description for each fine-grained persona. Additionally, we employ resisting strategies outlined by Dutt et al. (2021) to direct the behavior of simulators. Finally, our mixed-initiative role-play prompt for each agent includes the assigned persona, a set of resisting strategies, and conversation context. These elements aid in guiding user simulators to exhibit diverse non-collaborative behaviors. In total, we develop 300 diverse user simulators for each evaluation task, representing 20 persona categories (i.e., Big-Five Personality $\times$ Decision-Making Styles).
@@
Evaluation Tasks. In line with Deng et al. (2023b); Wang et al. (2019), we conduct experiments on two benchmark non-collaborative tasks: the price negotiation task, utilizing the test [^4] Our data split follows the previous study Deng et al. (2023c); Wang et al. (2019). dataset of CraigslistBargain (CB) He et al. (2018) and the charity persuasion task, employing the test dataset of PersuasionForGood (P4G) Wang et al. (2019). Notably, the dialogue agents play the role of buyer and persuader, respectively, to accomplish their goals.
@@
Evaluation Metrics. Following Deng et al. (2023c), we consider three commonly used metrics: Success Rate (SR), Average Turn (AT) and Sale-to-List Ratio (SL%). The SR measures effectiveness by the percentage of goal achievement within a maximum number of turns, while AT measures efficiency by the average number of turns required to achieve the goal. As for the CB task, we additionally adopt the SLZhou et al. (2019) to determine the effectiveness of goal completion. Formally, the SLis expressed as $(P_{deal}-P_{target}^{seller})/(P_{target}^{buyer}-P_{target}^{seller})$, where $P_{deal}$ is the final deal price, $P_{target}^{buyer}$ and $P_{target}^{seller}$ are the target prices of both parties. A higher SLrepresents the buyer gets more benefits from the deal. If failing to reach a deal at the end, we set SLas 0.
@@ 4.1 User-Aware Strategic Planning
Trip aims to explicitly infer user characteristics and then incorporate them into the strategic planning module, parameterized by a trainable BERT. In particular, building upon the advanced Theory-of-Mind capability of LLMs Sap et al. (2022); Moghaddam and Honey (2023), Trip captures users’ mental states and future possible actions during interactions to understand their interests and predicts how TRIP’s responses may influence them. In this case, mental states pertains to what they aim to accomplish, such as the target price or whether they will donate, while future actions relates to what the user is likely to discuss next Hu et al. (2023); Zhou et al. (2023a). Formally, given the dialogue history $D=(u_{1}^{sys},u_{1}^{usr},...,u_{t}^{sys},u_{t}^{usr})$, where $u_{i}^{sys}$ and $u_{i}^{usr}$ denote the $i$-th utterances of both parties and $t$ is the number of utterances, we feed the dialogue history $D$ into the LLM and prompt it to infer mental states $\mathcal{M}$ and future actions $\mathcal{F}$, i.e., $P_{LLM}(\mathcal{M},\mathcal{F}|D)$. Subsequently, we feed the $\mathcal{M},\mathcal{F},D$} into the strategy planner $\pi_{\theta}$ to predict the next strategy. The output space of $\pi_{\theta}$ is a set of strategies666e.g., the elicitation of specific emotions to influence other. pre-defined by Deng et al. (2023c); Wang et al. (2019), each of them is attached with a pre-defined natural language instructions.
@@ 4.2 Population-based Training Paradigm
Reward Design. Following Deng et al. (2023c), we prompt LLMs to judge the conversation progress at each turn and transform it into scalar rewards. Specifically, in the negotiation task, we employ a separate GPT3.5 OpenAI (2022) to assess whether both parties have reached a deal. In the persuasion task, we ask the GPT3.5-based user simulator to express its willingness to donation. Our rewards are determined based on three situations: 1) Successful goal achievement by the dialogue agent results in a significant positive reward, defined as 1.0 in the charity persuasion task and the value of SLin the price negotiation task. 2) Failure to achieve goals leads to a substantial negative reward of -1.0 for the dialogue agent. 3) Furthermore, we assign a small negative reward (-0.1) per turn to penalize the lengthy conversation, which promotes the efficient goal achievement.
@@ 5 Experiments
LLM-based baselines. We consider LLM-based dialogue agents with two types of strategic planning modules, as discussed in Section 2. 1) Prompt-based planning, including Standard, ProCoT Deng et al. (2023b) and ICL-AIF Fu et al. (2023), which use mixed-initiative prompts, CoT, and AI feedback to select next strategies, respectively. 2) External strategy planners, including GDP-MCTS Yu et al. (2023) and PPDPP Deng et al. (2023c), which utilize Monte Carlo Tree Search and a trainable plug-in for determining next-step strategies, respectively. Note that all baselines fail to model user-specific characteristics explicitly and are trained using one user simulator. Implementation details are presented in Appendix B.
@@ 5.1 Overall Performance
Trip is a promising method for achieving effective non-collaborative strategies tailored for diverse users. As illustrated in Table 2, Trip significantly outperforms all the baselines with a noticeable margin across two tasks. It not only efficiently achieves the conversational goal (less AT) but also effectively accomplishes tasks (higher SR and higher SL%). Moreover, as depicted in Figure 3, Trip shows balanced improvements across different user personas, significantly outperforming other agents by a substantial margin, in contrast to the biased improvements of PPDPP in Section 3.2. This suggests that Trip is capable of generating strategies that generalize well to diverse users. This also implies that the behavior pattern pf a single LLM-based user simulator is limited in scope. Moreover, our human evaluation results in Figure 4 show our Trip largely outperform the Standard and PPDPP when interacting with real users. This evidences the effectiveness and practical utility of our proposed Trip.
@@ 6 Conclusion
Sensitivity of Prompts. Similar to other studies on prompting LLMs Deng et al. (2023b), the evaluation results are expected to be influenced by the prompts. Following Deng et al. (2023c), we employ the mixed-initiative format to formulate our prompts, as it offers stability and control. The impact of prompts and their optimality present important areas of investigation within LLMs, calling for exploration in future studies.
Limited Non-collaborative Tasks. We only conduct our experiments on the two non-collaborative dialogue tasks (i.e., price negotiation and charity persuasion) due to their status as classic and widely-recognized benchmarks Deng et al. (2023b); Chawla et al. (2023a). In the future, we plan to apply our proposed Trip in a broader range of non-collaborative dialogue scenarios.

### v3

# Strength Lies in Differences! Improving Strategy Planning for Non-collaborative Dialogues via Diversified User Simulation

@@ Abstract
Strength Lies in Differences! Improving Strategy Planning for Non-collaborative Dialogues via Diversified User Simulation
@@ 1 Introduction
Non-collaborative dialogues, such as negotiation He et al. (2018) and persuasion Wang et al. (2019), occur when the agent and user hold conflicting interests Deng et al. (2023a, b); Lei et al. (2022). Typically, both parties need to employ various strategies to achieve an agreement favorable to themselves Keizer et al. (2017); Zhang et al. (2023b); Zhan et al. (2024). As user resistance varies depending on the agent’s strategies Shi et al. (2019); Dutt et al. (2021), it is imperative for the agent to perform strategic planning tailored to diverse users. Relying on a one-size-fits-all strategy can leave the agent vulnerable to others taking advantage due to its lack of adaptability and flexibility Yang et al. (2021); Deng et al. (2024); Xu et al. (2023).
Recent efforts have resorted to large language models (LLMs) as dialogue agents to perform non-collaborative tasks Deng et al. (2023d); Fu et al. (2023); Zhang et al. (2023a). They aim to guide the response of LLMs through mixed-initiative prompts Chen et al. (2023); Deng et al. (2023d); Zhang et al. (2023a) or incorporating an external strategy planner Yu et al. (2023); Deng et al. (2023e). However, these initiatives has been criticized regarding its performance in real-world scenarios Deng et al. (2023e); Kwon et al. (2024), where users have various non-collaborative strategies. We attribute this outcome to the neglect of two crucial aspects: 1) Existing methods fail to incorporate explicit user-specific characteristics into their strategic planning, instead relying solely on the conversational history. Importantly, by creating informative representations of individual users, agents can adapt their behaviors and devise tailored strategies Jang et al. (2020); Yang et al. (2021). 2) Their training paradigm fails to generate strategic planners that generalize well to diverse users. Their paradigms are oversimplified, relying on a single user simulator for interactive training. This simulator is restricted in generating varied non-collaborative behaviors, often exhibiting a focus on prioritizing user contentment Zhang et al. (2023c); Durmus et al. (2023); Bianchi et al. (2024). Essentially, agents trained in this manner are accustomed to engage with a single user exclusively, leading to rigidity and obstinacy when encountering new users with different interaction behaviors Wang et al. (2023); Safdari et al. (2023).
@@
To tackle these challenges, we design a simple yet effective method, called Trip, to improve LLMs’ capability in Tailored stRategIc Planning. Trip includes a user-aware strategic planning module and a population-based training paradigm. Specifically, the strategic planning module incorporates user-specific characteristics into strategic planning using the Theory-of-Mind (ToM) Premack and Woodruff (1978); Wimmer and Perner (1983). This involves analyzing users’ mental states and future possible actions during interactions to understand their interests Yang et al. (2021); Chawla et al. (2023a). Moreover, instead of relying on a solitary user simulator, our population-based training paradigm promotes the adaptation of the strategic planning module to various users, achieved by training it with more diverse user simulators. Each simulator is equipped with extensive sets of non-collaborative strategies and role-playing personas Chen et al. (2024). As such, Trip essentially manipulates the experience of the dialogue agent, enabling it to recognize the importance of tailoring strategies for individual users. Our key contributions are concluded below:
@@ 2 Related Work
Strategic planning for non-collaborative dialogues. Recent researches have introduced various methods based on LLMs to enhance their effectiveness in strategic planning. These methods can be categorized into two types: 1) Developing stimulus prompts to unleash the potential of LLMs. Chen et al. (2023) validate the effectiveness of using mixed-initiative prompts to tackle proactive dialogue challenges. Deng et al. (2023d) and Zhang et al. (2023a) encourage LLMs to engage in self-reflection to plan their next actions. Fu et al. (2023) employ self-play simulations to iteratively refine strategic planning by soliciting feedback from other LLMs. Nonetheless, as highlighted by Deng et al. (2023e), the effectiveness of these approaches is impeded by non-trainable parameters. 2) Equipping LLMs with an external strategy planner. The planner is capable of generating prompts at each turn, providing nuanced, instance-specific guidance and control over LLMs. This could be integrated using methods like Monte Carlo Tree Search Yu et al. (2023) or a plug-in model Deng et al. (2023e), which can be fine-tuned for improving the strategic planning capability without affecting the functionalities of LLM-powered dialogue agents. However, these methods still struggle to achieve promising results due to their inability to integrate user-specific characteristics into their strategic planning. Complementary to Deng et al. (2023e), our work investigates the importance of tailored strategic planning by modeling user-related characteristics explicitly.
Training paradigms for non-collaborative dialogues. Current training paradigms involve the dialogue agent interacting with a single user simulator to enhance its strategic planning capabilities. In specific, Chawla et al. (2023b) build a user simulator that mimics human-human dialogue data in a supervised manner, while Yu et al. (2023); Deng et al. (2023e) resort to a role-playing LLM-based user simulator. However, a single user simulator can only represent the behaviors of one or a type of users, potentially leading to the under-representation of other users’ behaviors, as evidenced by Liu et al. (2023); Shi et al. (2019). Therefore, existing training paradigms fail to produce strategic planners that cater to diverse users with varying behaviors. In this paper, our work investigates the importance of tailored strategic planning by diversifying the user’s behaviors using population-based training.
@@ 3.1 Evaluation Setup
Evaluation Overview. The environment encompasses various synthetic user simulators showcasing diverse non-collaborative behaviors. In the evaluation process, each dialogue agent must interact with these simulators Deng et al. (2023e). During their interactions, the dialogue agent and user simulator alternate in employing strategies in their responses with the ultimate aim of maximizing their own self-interest. The interactions continues until the conversational goal is achieved or the maximum number of turns is reached. We gather these interactions and assess the agents performances.
@@
Baselines. We consider two representative baselines: Standard agent (i.e., vanilla LLM without any modification) and PPDPP agent Deng et al. (2023e), which is current SOTA agent with a trainable external strategy planner [^1] Notably, we also consider other existing dialogue agents in our main experiments..
@@
Diverse User Simulators. Our simulators are synthesized with non-collaborative behaviors, guided by their task-relevant personas. As evidenced by previous study Deng et al. (2023c); Bianchi et al. (2024); Huang et al. (2024), LLMs are limited to demonstrate non-collaborative behaviors. To this end, we prompt non-collaborative behaviors explicitly into LLMs using the resisting strategies that are designed to foil persuasion attempts Fransen et al. (2015); Tian et al. (2020); Dutt et al. (2021). Initially, we equip LLMs with different personas Jiang et al. (2023); Zhou et al. (2023b); Zhang et al. (2023b), which are used to select non-collaborative behaviors from the set of resisting strategies. Following Wang et al. (2019); Jiang et al. (2024), we consider two types of personas, including Big-Five Personality [^2] Openness, Conscientiousness, Extraversion, Agreeableness, and Neuroticism Goldberg (1992) and Decision-Making Styles [^3] Directive, Conceptual, Analytical, and Behavioral Scott and Bruce (1995), together with LLM-generated cohesive description for each fine-grained persona. Additionally, we employ resisting strategies outlined by Dutt et al. (2021) to direct the behavior of simulators. Finally, our mixed-initiative role-play prompt for each agent includes the assigned persona, a set of resisting strategies, and conversation context. These elements aid in guiding user simulators to exhibit diverse non-collaborative behaviors. In total, we develop 300 diverse user simulators for each evaluation task, representing 20 persona categories (i.e., Big-Five Personality $\times$ Decision-Making Styles).
@@
Evaluation Tasks. In line with Deng et al. (2023d); Wang et al. (2019), we conduct experiments on two benchmark non-collaborative tasks: the price negotiation task, utilizing the test [^4] Our data split follows the previous study Deng et al. (2023e); Wang et al. (2019). dataset of CraigslistBargain (CB) He et al. (2018) and the charity persuasion task, employing the test dataset of PersuasionForGood (P4G) Wang et al. (2019). Notably, the dialogue agents play the role of buyer and persuader, respectively, to accomplish their goals.
@@
Evaluation Metrics. Following Deng et al. (2023e), we consider three commonly used metrics: Success Rate (SR), Average Turn (AT) and Sale-to-List Ratio (SL%). The SR measures effectiveness by the percentage of goal achievement within a maximum number of turns, while AT measures efficiency by the average number of turns required to achieve the goal. As for the CB task, we additionally adopt the SLZhou et al. (2019) to determine the effectiveness of goal completion. Formally, the SLis expressed as $(P_{deal}-P_{target}^{seller})/(P_{target}^{buyer}-P_{target}^{seller})$, where $P_{deal}$ is the final deal price, $P_{target}^{buyer}$ and $P_{target}^{seller}$ are the target prices of both parties. A higher SLrepresents the buyer gets more benefits from the deal. If failing to reach a deal at the end, we set SLas 0.
@@ 4.1 User-Aware Strategic Planning
Trip aims to explicitly infer user characteristics and then incorporate them into the strategic planning module, parameterized by a trainable BERT. In particular, building upon the advanced Theory-of-Mind capability of LLMs Sap et al. (2022); Moghaddam and Honey (2023), Trip captures users’ mental states and future possible actions during interactions to understand their interests and predicts how TRIP’s responses may influence them. In this case, mental states pertains to what they aim to accomplish, such as the target price or whether they will donate, while future actions relates to what the user is likely to discuss next Hu et al. (2023); Zhou et al. (2023a). Formally, given the dialogue history $D=(u_{1}^{sys},u_{1}^{usr},...,u_{t}^{sys},u_{t}^{usr})$, where $u_{i}^{sys}$ and $u_{i}^{usr}$ denote the $i$-th utterances of both parties and $t$ is the number of utterances, we feed the dialogue history $D$ into the LLM and prompt it to infer mental states $\mathcal{M}$ and future actions $\mathcal{F}$ in an open-ended manner, i.e., $P_{LLM}(\mathcal{M},\mathcal{F}|D)$. Subsequently, we feed the $\mathcal{M},\mathcal{F},D$} into the strategy planner $\pi_{\theta}$ to predict the next strategy. The output space of $\pi_{\theta}$ is a set of strategies666e.g., the elicitation of specific emotions to influence other. pre-defined by Deng et al. (2023e); Wang et al. (2019), each of them is attached with a pre-defined natural language instructions.
@@ 4.2 Population-based Training Paradigm
Reward Design. Following Deng et al. (2023e), we prompt LLMs to judge the conversation progress at each turn and transform it into scalar rewards. Specifically, in the negotiation task, we employ a separate GPT3.5 OpenAI (2022) to assess whether both parties have reached a deal. In the persuasion task, we ask the GPT3.5-based user simulator to express its willingness to donation. Our rewards are determined based on three situations: 1) Successful goal achievement by the dialogue agent results in a significant positive reward, defined as 1.0 in the charity persuasion task and the value of SLin the price negotiation task. 2) Failure to achieve goals leads to a substantial negative reward of -1.0 for the dialogue agent. 3) Furthermore, we assign a small negative reward (-0.1) per turn to penalize the lengthy conversation, which promotes the efficient goal achievement.
@@ 5 Experiments
LLM-based baselines. We consider LLM-based dialogue agents with two types of strategic planning modules, as discussed in Section 2. 1) Prompt-based planning, including Standard, ProCoT Deng et al. (2023d) and ICL-AIF Fu et al. (2023), which use mixed-initiative prompts, CoT, and AI feedback to select next strategies, respectively. 2) External strategy planners, including GDP-MCTS Yu et al. (2023) and PPDPP Deng et al. (2023e), which utilize Monte Carlo Tree Search and a trainable plug-in for determining next-step strategies, respectively. Note that all baselines fail to model user-specific characteristics explicitly and are trained using one user simulator. Implementation details are presented in Appendix B.
@@ 5.1 Overall Performance
Trip is a promising method for achieving effective non-collaborative strategies tailored for diverse users. As illustrated in Table 2, Trip significantly outperforms all the baselines with a noticeable margin across two tasks. It not only efficiently achieves the conversational goal (less AT) but also effectively accomplishes tasks (higher SR and higher SL%). Moreover, as depicted in Figure 3, Trip shows balanced improvements across different user personas, significantly outperforming other agents by a substantial margin, in contrast to the biased improvements of PPDPP in Section 3.2. This suggests that Trip is capable of generating strategies that generalize well to diverse users. This also implies that the behavior pattern pf a single LLM-based user simulator is limited in scope. Moreover, our human evaluation results in Figure 4 show our Trip largely outperform the Standard and PPDPP when interacting with real users. Notably, we observed that PPDPP does not consistently surpass the Standard approach across the two tasks. For instance, while it achieves a higher success rate in the negotiation task, it necessitates more interaction rounds. This evidences the effectiveness and practical utility of our proposed Trip.
@@ 6 Conclusion
Sensitivity of Prompts. Similar to other studies on prompting LLMs Deng et al. (2023d), the evaluation results are expected to be influenced by the prompts. Following Deng et al. (2023e), we employ the mixed-initiative format to formulate our prompts, as it offers stability and control. The impact of prompts and their optimality present important areas of investigation within LLMs, calling for exploration in future studies.
Limited Non-collaborative Tasks. We only conduct our experiments on the two non-collaborative dialogue tasks (i.e., price negotiation and charity persuasion) due to their status as classic and widely-recognized benchmarks Deng et al. (2023d); Chawla et al. (2023a). In the future, we plan to apply our proposed Trip in a broader range of non-collaborative dialogue scenarios Zhang et al. (2024); Zhou et al. (2023b).
Acknowledgements
This work was supported in part by the National Natural Science Foundation of China (No. 62272330 and No. 62206191); in part by the Natural Science Foundation of Sichuan (No. 2023NSFSC0473); in part by the Fundamental Research Funds for the Central Universities (No. 2023SCU12089 and No. YJ202219); in part by the Singapore Ministry of Education (MOE) Academic Research Fund (AcRF) Tier 1 grant (No. MSS24C004).